The slope of the AE curve is .80. What is the multiplier? Everything else the same, by how much does equilibrium aggregate expenditure increase if

a)  exports increase from $1.75 trillion to $2.25 trillion.                             (8 marks)

b)  government expenditure on goods and services decrease from $2.0 trillion to $1.8 trillion. (6 marks)

c)  investment increases from $1.2 trillion to $2.3 trillion.                         (6 marks)



1
Expert's answer
2021-09-20T11:58:20-0400

Slope of AE = MPC = 0.8

Expenditure multiplier = "\frac{1}{1-mpc}" = "\frac{1}{1-0.8}" = 5


Therefore expenditure multiplier is 0.5


a. Change in exports = 2.25 - 1.75 = 0.5 trillion

As per the multiplier formula,

"\frac{change in AE}{change in exports}" = 5

Therefore, change in AE = Change in exports "\times" 5 = (0.5)"\times" 5 = $2.5 trillion

Therefore, AE will increase by $2.5 trillion.

 

b. Change in G = 1.8 - 2 = -0.2 trillion

According to the multiplier formula,

"\frac{Change in AE}{change in G} = 5"

Therefore, change in AE = Change in G "\times" 5 = (-0.2) "\times" 5 = $-1 trillion

Therefore, AE will decrease by $1 trillion.

 

c. Change in I = 2.3 - 1.2 = 1.1 trillion

According to the multiplier formula,

"\frac{change in AE}{change in I} = 5"

Therefore, change in AE = Change in I"\times" 5 = (1.1)"\times" 5 = $5.5 trillion

Therefore, AE will increase by $5.5 trillion.